OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10

OpenStudy (anonymous):

x = 0 x = −3 and x = 1 x = −1 and x = 3 No solution

OpenStudy (anonymous):

@tanya123

OpenStudy (anonymous):

@hba @jim_thompson5910

jimthompson5910 (jim_thompson5910):

First you must isolate the absolute value. So divide both sides by -5 to get \[\Large -5|x+1| = 10\] \[\Large \frac{-5|x+1|}{-5} = \frac{10}{-5}\] \[\Large \frac{\cancel{-5}|x+1|}{\cancel{-5}} = \frac{10}{-5}\] \[\Large |x+1| = -2\] What do you notice about the last equation?

OpenStudy (anonymous):

The absolute value is by itself?

jimthompson5910 (jim_thompson5910):

what else

OpenStudy (anonymous):

I don't know

jimthompson5910 (jim_thompson5910):

what does absolute value represent in general?

OpenStudy (anonymous):

The value of a number without signs or anything

jimthompson5910 (jim_thompson5910):

it also represents distance for example |-7| = 7 and this says "the number -7 is seven units away from 0"

jimthompson5910 (jim_thompson5910):

since negative distance makes no sense, it means that the result of an absolute value expression is never negative

jimthompson5910 (jim_thompson5910):

Saying \[\Large |x+1| = -2\] makes no sense. So therefore, there are no solutions to that equation.

jimthompson5910 (jim_thompson5910):

There is no way to make the left side of \[\Large |x+1| = -2\] equal to -2 (or any negative number for that matter) regardless of any x value you pick.

OpenStudy (anonymous):

So if the answer is a negative number, there is no solution?

jimthompson5910 (jim_thompson5910):

if you have something of the form |x| = -2 or |x| = -7, then it's not possible so there are no solutions

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

np